# DEL-Ranking: Ranking-Correction Denoising Framework for Elucidating Molecular Affinities in DNA-Encoded Libraries

### v1

@@ Abstract
DNA-encoded library (DEL) screening has revolutionized the detection of protein-ligand interactions through read counts, enabling the rapid exploration of vast chemical spaces. However, noise in read counts, stemming from nonspecific interactions, can mislead this exploration process. Neural networks trained on DEL libraries have been employed to discern and capture nuanced, task-specific patterns within chemical landscapes. However, existing methods overlook two critical aspects: (1) the inherent ranking nature of read counts and (2) the potential of true activity labels to correct systematic biases. We present DEL-Ranking, a novel distribution-correction denoising framework that addresses these challenges. Our approach introduces two key innovations: (1) a novel ranking loss that rectifies relative magnitude relationships between read counts enabling the learning of causal features determining activity levels, and (2) an iterative algorithm employing self-training and consistency loss to establish model coherence between activity label and read count predictions. Furthermore, we contribute three new DEL screening datasets, the first to comprehensively include multi-dimensional molecular representations, protein-ligand enrichment values, and their activity labels. These datasets mitigate data scarcity and incomplete data issues in AI-driven DEL screening research, providing novel benchmark datasets for this field. Rigorous evaluation on diverse DEL datasets demonstrates DEL-Ranking’s superior performance across multiple correlation metrics, with significant improvements in binding affinity prediction accuracy. Our model exhibits zero-shot generalization ability across different protein targets and successfully identifies potential motifs determining compound binding affinity. Our work not only advances the field of DEL screening analysis but also provides valuable resources for future research in this area.
@@ 1 Introduction
DNA-encoded library (DEL) technology has emerged as a revolutionary approach for protein-ligand binding detection, offering unprecedented advantages over traditional high-throughput screening methods(Franzini et al., 2014; Neri & Lerner, 2018; Peterson & Liu, 2023; Ma et al., 2023). The DEL screening process involves multiple stages including cycling, binding, washing, elution, and amplification (as shown in Figure 1). This process generates large-scale read count data, serving as a proxy for potential binding affinity(Machutta et al., 2017; Foley et al., 2021). The read counts represent the frequency of each compound in the selected pool after undergoing target protein binding and subsequent processing steps (Favalli et al., 2018). These read counts typically include control counts and protein counts, representing values corresponding to the scenarios without and with specific protein targets, respectively.
DEL screening enables the rapid and cost-effective evaluation of vast chemical spaces, typically encompassing billions of compounds (Satz et al., 2022), against biological targets of interest (Neri & Lerner, 2017). This approach has gained widespread adoption in drug discovery due to its ability to identify novel chemical scaffolds and accelerate lead compound identification (Brenner & Lerner, 1992; Goodnow Jr & Davie, 2017; Yuen & Franzini, 2017). This study aims to enhance the correlation between predicted read counts and true binding affinity. However, DEL screening read counts are subject to $Distribution~{}Noise$, stemming from both experimental variabilities and intrinsic factors (Favalli et al., 2018). These noisy elements introduce discrepancies between read counts and experimentally determined $K_{i}$ values (Yung-Chi & Prusoff, 1973), which represent the binding affinity of a ligand to its target protein in the training data, making this task highly challenging.
@@
To mitigate noise, researchers apply threshold-based filtering to enrichment factors, calculated as the ratio of protein count to control count(Gu et al., 2008; Kuai et al., 2018), to classify compounds as potential hits or noise(Gu et al., 2008; Kuai et al., 2018). Although these methods offer advantages in interpretability and computational efficiency, they focus solely on the inherent properties of read counts, overlooking the complex relationships between ligand molecular structures and their corresponding read counts(McCloskey et al., 2020). To address this limitation, researchers leveraged machine learning techniques to capture the non-linear relationships between ligand molecules and their count labels, aiming to predict more accurate read counts (McCloskey et al., 2020; Ma et al., 2021). These approaches, while more sophisticated in modeling molecular features, initially did not consider the constraints at the read count distribution level. Recognizing this gap, researchers further advanced the field by unifying both molecular structure-based predictions and distribution-level constraints. They developed methods incorporating distribution consistency losses, which use ligand sequence embedding to predict enrichment factors and evaluate the consistency of protein and control counts (Lim et al., 2022; Hou et al., 2023).
Recognizing the limitations of 2D ligand sequences in capturing spatial information, DEL-Dock (Shmilovich et al., 2023) introduced protein-ligand 3D conformational information to enhance denoising. However, compared to exploring different dimensional representations of ligands, enrichment factors only focus on the absolute values of ligand-corresponding read counts, neglecting their relative ordering, which is more stable and reliable, especially with experimental noise and systematic errors. Moreover, read counts primarily reflect the enrichment degree of compounds in the screening process but do not directly represent the biological activity of compounds. Without incorporating the activity information, models may overlook important ligand-activity relationships.
To address these multifaceted challenges, we propose a novel denoising framework that integrates a theoretically grounded combined ranking loss with an iterative validation loop for read count distribution correction. We introduce two novel constraints: the Pair-wise Soft Rank (PSR) and List-wise Global Rank (LGR). These constraints enable simultaneous learning of local discriminative features and global patterns in read count distributions. By emphasizing the relative relationships between compounds while integrating point-wise constraints based on their absolute read count values, our method provides a comprehensive approach to modeling the complex distributions in DEL. We prove that our novel ranking loss provides complementary information to ZIP loss and reduces the expected error bound, thereby ensuring the reliability of our algorithm.
Considering the ligand-activity relationships for distribution correction, we introduce a novel activity label derived from the analysis of functional groups in the DEL library’s ligands. This innovation forms the foundation of our proposed Activity-Referenced Distribution Correction (ARDC) framework. ARDC is designed to optimize the correlation between predicted read count values and potential binding affinities. It incorporates two key components: Iterative Activity-Guided Refinement (AGR) and Consistency-Driven Error Correction (CDEC). The AGR process, inspired by self-training techniques (Zoph et al., 2020), iteratively refines predictions by adjusting the model based on evolving read count to binding affinity relationships. CDEC addresses potential error accumulation through a consistency loss function, aligning predictions with biological reality and resolving discrepancies between the read counts and the activity levels.
To address the information gap in existing open-source datasets, we introduce three comprehensive datasets incorporating ligand 2D sequences, 3D structures, and corresponding activity labels. These datasets establish new benchmarks for denoising DEL libraries. To validate our approach, we conducted extensive experiments on five diverse DEL datasets. Our results demonstrate state-of-the-art performance across various correlation metrics, with comprehensive ablation studies underscoring the superiority of our distribution correction method. This work represents a significant advancement towards more accurate and reliable binding affinity predictions from DEL screening data, potentially accelerating drug discovery and enhancing the identification of potent lead compounds.
@@ 2 Related Works
While existing methods offer improved scalability and the ability to capture complex molecular interactions, they still face challenges in interpretability. Moreover, they often lack a robust theoretical foundation for handling the unique characteristics of DEL read count data, particularly in terms of ranking and distribution correction. The persistent issues of Distribution Noise and Distribution Shift in DEL data remain inadequately addressed by current approaches.
@@ 3 Method
Figure 2: Overview of DEL-Ranking framework. The model directly fuses molecule binding poses and fingerprints as input features. ARDC employs target effects and binding affinity to enhance read count prediction. The ranking-based loss incorporates target effects and matrix effects for noise removal, improving the correlation between predicted read counts and true binding affinities.
@@ 3.1 Problem Formulation and preliminaries
DEL Prediction Framework. Given a DEL dataset $\mathcal{D}=\{(\mathbf{f}_{i},\mathbf{p}_{i},M_{i},R_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{f}_{i}\in\mathbb{R}^{d}$ denotes the molecular fingerprint, $\mathbf{p}_{i}\in\mathbb{R}^{m}$ represents the binding pose, $M_{i}\in\mathbb{R}^{+}$ is the matrix count derived from control experiments without protein targets, $R_{i}\in\mathbb{R}^{+}$ is the target count obtained from experiments involving protein target binding, and $y_{i}\in\{0,1\}$ indicates the activity label. We propose a joint multi-task learning framework $\mathcal{F}:\mathbb{R}^{d}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{+}\times\mathbb{R}^{+}\times[0,1]$ such that:
@@
Zero-Inflated Distribution. DEL screening often results in read count distributions with a high proportion of zeros due to experimental factors. To address this, previous methods (Shmilovich et al., 2023; Lim et al., 2022) have employed zero-inflated distributions, modeling read counts $r_{i}$ as ($M_{i}$, $R_{i}$), where $M_{i}$ accounts for excess zeros and $R_{i}$ represents non-zero counts.
$P(r_{i}|\pi_{i},\mu_{i},\phi)=\begin{cases}\pi_{i}+(1-\pi_{i})(1+\phi\mu_{i})^{-1/\phi},&\text{if }r_{i}=0\\ (1-\pi_{i})\frac{\Gamma(r_{i}+1/\phi)}{r_{i}!\Gamma(1/\phi)}\left(\frac{1}{1+\phi\mu_{i}}\right)^{1/\phi}\left(\frac{\phi\mu_{i}}{1+\phi\mu_{i}}\right)^{r_{i}},&\text{if }r_{i}>0\end{cases}$ (2)
where $\pi_{i}$ denotes the probability of excess zeros, $\mu_{i}$ represents the mean of the negative binomial distribution, and $\phi$ is the dispersion parameter.
$K_{i}$ Estimation. DEL read count prediction ultimately aims to estimate compound-target binding affinities ($K_{i}$ values), crucial for identifying promising drug candidates. We evaluate our model’s effectiveness using the Spearman rank correlation coefficient ($\rho_{s}$) between predicted read counts and true $K_{i}$ values: $\rho_{s}=1-\frac{6\sum d^{2}}{n(n^{2}-1)}$ where $n$ is the sample number and $d$ is the ranking discrepancy between predictions and true values. Ideally, $K_{i}$ values and read counts are negatively correlated.
@@ 3.2 Ranking-based Distribution Noise Removal
where $\beta\in[0,1]$ is a balancing hyperparameter. $\mathcal{L}_{\text{PSR}}$ (Pairwise-Soft Ranking Loss) addresses local pairwise comparisons, while $\mathcal{L}_{\text{LGR}}$ (listwise Global Ranking Loss) captures global ranking information. Together, they aim to achieve a well-ordered Zero-Inflated Poisson distribution for read count values. To formally establish the effectiveness of our ranking-based approach, we provide the following theoretical justification:
@@ 3.2.1 Pairwise Soft Ranking Loss
To address the challenges of modeling relative discrete connections between compound pairs and remove read count Distribution Noise, we propose the differentiable Pairwise Soft Ranking loss $\mathcal{L}_{\text{PSR}}$. This novel loss function is designed to capture fine-grained ranking information while maintaining smooth gradients for stable optimization. $\mathcal{L}_{\text{PSR}}$ is expressed as follows:
$\mathcal{L}_{PSR}(y_{i},y_{j},T)$ $=-\sum_{i=1}^{n}\hat{y}_{i}\left(\sum_{j\neq i}(\Delta_{ij}\cdot\sigma_{ij})-\sum_{j\neq i}(\Delta_{ji}\cdot\sigma_{ji})\right)$ (4)
$\sigma_{ij}$ $=\frac{1}{1+e^{-|\hat{y}_{i}-\hat{y}_{j}|/T}},\quad\Delta_{ij}=\frac{\Delta\text{G}_{ij}\cdot\Delta\text{D}_{ij}}{Z}$
where $\hat{y}_{i}$ represents the predicted read count value for compound $i$. To ensure smooth gradients and numerical stability, we introduce a scaling factor $\sigma_{ij}$ with temperature $T$, allowing for fine-tuned control over the ranking sensitivity.
To accurately capture the impact of ranking changes, we design $\Delta_{ij}$ to approximate the relative importance of swapping compounds $i$ and $j$. We introduce a gain function $G_{i}=\text{softplus}(y_{i})$ and a rank-based discount function $D_{i}=1/(\log_{2}(\text{rank}_{i}+1)+\epsilon)$, where $\epsilon$ is a small constant for numerical stability. These smooth functions ensure stable gradients during optimization.
To further normalize the impact of ranking changes, we compute a normalization factor in $\Delta\text{D}_{ij}$:
$Z=\sum_{k=1}^{K}\frac{\text{softplus}(y_{[k]})}{\log_{2}(k+1)+\epsilon}$ (5)
where $y_{[k]}$ is the k-th element of the descending-sorted y. This factor adapts our loss to varying dataset sizes and read count distributions, enhancing the robustness of our ranking model.
@@
$\mathcal{L}_{\text{LGR}}(s_{i},s_{j},T)=-\sum_{i=1}^{n}\log\frac{\exp(s_{\pi(i)}/{T})}{\sum_{j=i}^{n}\exp(s_{\pi(i)}/{T})}+\sigma\mathcal{L}_{\text{con}}(s_{i},s_{j})$ (6)
where $\pi$ is the true ranking permutation of the compounds; $T$ is a temperature parameter for score rescaling to sharpen the predicted distribution, and $\mathcal{L}_{\text{con}}$ denotes a contrastive loss among ranking scores to capture local connections, and $\sigma$ denotes the weight. This formulation is designed to achieve two critical objectives in DEL experiments: (1) Near-deterministic selection of compounds with the highest read counts, corresponding to the highest binding affinities; (2) Increased robustness to small noise perturbations in the experimental data. As $T$ approaches 0, our model becomes increasingly selective towards high-affinity compounds while maintaining resilience against common experimental noises. This dual optimization leads to more consistent identification of promising drug candidates and enhanced reliability in the face of experimental variability.
Despite the strengths of $\mathcal{L}_{\text{PSR}}$ and $\mathcal{L}_{\text{LGR}}$, they struggle to differentiate activity levels among compounds with identical read count values, particularly those affected by experimental noise. This limitation can lead to misclassification of high-activity samples with artificially low read counts as truly low-activity samples. To address this critical issue, we introduce a novel contrastive loss function $\mathcal{L}_{\text{con}}$, designed to enhance discrimination between varying levels of biological activity, especially for samples with zero or identical read count values. Let $f:\mathcal{S}\rightarrow\mathbb{R}$ be a ranking function and $\tau>0$ a fixed threshold. We define $\mathcal{L}_{\text{con}}:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}{\geq 0}$ as:
$\mathcal{L}_{\text{con}}(s_{i},s_{j})=\max\{0,\tau-(f(s_{i})-f(s_{j}))\}$ (7)
This loss function is positive if and only if $f(s_{i})-f(s_{j})<\tau$, enforcing a minimum margin $\tau$ between differently ranked samples. The constant gradients $\partial\mathcal{L}_{\text{con}}/\partial f(s_{i})=-1$ and $\partial\mathcal{L}_{\text{con}}/\partial f(s_{j})=1$ for $f(s_{i})-f(s_{j})<\tau$ promote robust ranking relationships.
Algorithm 1 Activity-Guided Refinement
1:Pose embeddings $\mathbf{h}_{p}$, Fingerprint embeddings $\mathbf{h}_{f}$, num_iterations $\mathbf{n}$, use_feedback
2:$x\leftarrow\texttt{PostAddLayer}(\mathbf{h}_{p}+\mathbf{h}_{f})$
3:$M_{\text{pred}}\leftarrow\texttt{MatrixHead}(\mathbf{h}_{f})$
4:Initialize $R_{\text{pred}}\leftarrow\mathbf{0}$, $p_{\text{pred}}\leftarrow\mathbf{0}$
@@
7:         $R_{\text{input}}\leftarrow[x;p_{\text{pred}}]$, $p_{\text{input}}\leftarrow[x;R_{\text{pred}}]$
@@
9:         $R_{\text{input}}\leftarrow x$, $p_{\text{input}}\leftarrow x$
@@
11:     $R_{\text{pred}}\leftarrow\texttt{EnrichmentHead}(\texttt{ReadHead}(R_{\text{input}}))$
12:     $p_{\text{pred}}\leftarrow\texttt{ActHead}(p_{\text{input}})$
13:end forreturn $M_{\text{pred}}$, $R_{\text{pred}}$, $p_{\text{pred}}$

### v2

@@ Abstract
DNA-encoded library (DEL) screening has revolutionized protein-ligand binding detection, enabling rapid exploration of vast chemical spaces through read count analysis. However, two critical challenges limit its effectiveness: distribution noise in low copy number regimes and systematic shifts between read counts and true binding affinities. We present DEL-Ranking, a comprehensive framework that simultaneously addresses both challenges through innovative ranking-based denoising and activity-referenced correction. Our approach introduces a dual-perspective ranking strategy combining Pair-wise Soft Rank (PSR) and List-wise Global Rank (LGR) constraints to preserve both local and global count relationships. Additionally, we develop an Activity-Referenced Correction (ARC) module that bridges the gap between read counts and binding affinities through iterative refinement and biological consistency enforcement. Another key contribution of this work is the curation and release of three comprehensive DEL datasets that uniquely combine ligand 2D sequences, 3D conformational information, and experimentally validated activity labels. We validate our framework on five diverse DEL datasets and introduce three new comprehensive datasets featuring 2D sequences, 3D structures, and activity labels. DEL-Ranking achieves state-of-the-art performance across multiple correlation metrics and demonstrates strong generalization ability across different protein targets. Importantly, our approach successfully identifies key functional groups associated with binding affinity, providing actionable insights for drug discovery. This work advances both the accuracy and interpretability of DEL screening, while contributing valuable datasets for future research.
@@ 1 Introduction
DNA-encoded library (DEL) technology has emerged as a revolutionary approach for protein-ligand binding detection, offering unprecedented advantages over traditional high-throughput screening methods(Franzini et al., 2014; Neri & Lerner, 2018; Peterson & Liu, 2023; Ma et al., 2023). The DEL screening process involves multiple stages including cycling, binding, washing, elution, and amplification (as shown in Figure 1). This process generates large-scale read count data, serving as a proxy for potential binding affinity(Machutta et al., 2017; Foley et al., 2021). The read counts represent the frequency of each compound in the selected pool after undergoing target protein binding and subsequent processing steps (Favalli et al., 2018). These read counts typically include matrix counts and target counts, representing values corresponding to the scenarios without and with specific protein targets, respectively.
DEL screening enables rapid and cost-effective evaluation of vast chemical spaces, typically encompassing billions of compounds (Satz et al., 2022), against biological targets (Neri & Lerner, 2017). This approach has gained widespread adoption in drug discovery due to its ability to identify novel chemical scaffolds and accelerate lead compound identification (Brenner & Lerner, 1992; Goodnow Jr & Davie, 2017; Yuen & Franzini, 2017). Despite its advantages, DEL screening faces two critical challenges: (1) DEL screening read counts are subject to Distribution Noise, stemming from both experimental factors and intrinsic library characteristics. This noise predominantly affects the low copy number regime ($<$10 copies) where read counts follow a Poisson distribution, creating significant discrepancies between observed counts and true binding affinities (Ki values) (Kuai et al., 2018; Favalli et al., 2018); and (2) DEL screening data faces Distribution Shift, where the mapping from read counts to binding affinities is systematically biased. While read counts reflect initial binding events, true binding affinity (Ki value) depends on multiple molecular properties that cannot be captured by enrichment data alone, creating a fundamental gap between read count-based and actual affinity distributions (Yung-Chi & Prusoff, 1973; Kuai et al., 2018). This study aims to address both challenges by enhancing the correlation between predicted read counts and true binding affinity.
@@
To address these challenges, various approaches have been developed. Early methods focused on mitigating Distribution Noise through threshold-based filtering of enrichment factors calculated as target count to matrix count ratios (Gu et al., 2008; Kuai et al., 2018). While computationally efficient and interpretable, these methods only consider read count properties, ignoring the complex relationships between molecular structures and their corresponding counts (McCloskey et al., 2020). Recent advances have taken two main directions. First, machine learning approaches were introduced to capture non-linear relationships between ligand molecules and their count labels (McCloskey et al., 2020; Ma et al., 2021). This was further enhanced by incorporating distribution-level constraints, using ligand sequence embeddings to jointly predict enrichment factors and ensure count consistency (Lim et al., 2022; Hou et al., 2023). Second, recognizing the limitations of 2D representations, DEL-Dock (Shmilovich et al., 2023) introduced 3D conformational information to improve denoising. By incorporating Zero-Inflated Poisson distribution (ZIP) modeling with 3D structural information, DEL-Dock achieved superior performance in read count prediction and demonstrated strong generalization ability across different protein targets.
A critical yet often overlooked aspect in current approaches is the biological activity information inherent in molecular structures. While the presence of certain functional groups, such as benzene sulfonamide, can serve as reliable indicators of binding potential and provide complementary information to read counts (Hou et al., 2023; leash-BELKA), and this prior knowledge can be extended to other DEL targets and datasets, current methods still face significant limitations in fully utilizing this information (wichert2024challenges). These methods primarily focus on absolute read count values instead of more robust relative ordering information, and rely solely on enrichment data without incorporating crucial biological activity labels. Although recent advanced computational approaches have made progress in addressing Distribution Noise, the fundamental challenge of Distribution Shift between enrichment data and true binding affinities remains largely unaddressed in current literature. This gap underscores the need for methods that can better integrate both functional group activity indicators and binding affinity information to improve both read count regression and the discovery of novel high-affinity functional groups.
To address these multifaceted challenges, we propose DEL-Ranking, an innovative framework that synergistically tackles both Distribution Noise and Distribution Shift. Our approach uniquely combines theoretically grounded ranking constraints with activity-referenced correction mechanisms that leverage molecular structural features, enabling robust read count denoising while preserving biological relevance. At the core of DEL-Ranking’s Distribution Noise mitigation is a novel dual-perspective ranking strategy. We introduce two complementary constraints: the Pair-wise Soft Rank (PSR) that captures local discriminative features, and the List-wise Global Rank (LGR) that maintains global distribution patterns. This ranking-based approach emphasizes relative relationships while preserving absolute count information, theoretically complementing the existing ZIP-based regression loss and reducing the expected error bound. To address Distribution Shift, we develop the Activity-Referenced Correction (ARC) module that bridges the gap between read counts and binding affinities. ARC operates through a two-stage process: a Refinement Stage that leverages self-training techniques (Zoph et al., 2020) for iterative optimization, and a Correction Stage that employs targeted consistency loss to ensure biological relevance. By incorporating novel activity labels derived from ligand functional group analysis, ARC effectively aligns read count predictions with true binding properties.
We validate our approach through extensive experiments on five diverse DEL datasets and introduce three new comprehensive datasets featuring ligand 2D sequences, 3D structures, and activity labels. Our method not only achieves state-of-the-art performance across multiple metrics but also successfully identifies key functional groups associated with high binding affinities. Through detailed ablation studies, we demonstrate each component’s effectiveness and show how our activity-referenced framework can reveal structure-activity relationships that guide rational drug design. This work advances both DEL screening accuracy and interpretability, potentially accelerating drug discovery through more precise binding affinity predictions and actionable structural insights.
@@ 2 Related Works
Existing methods demonstrate improved scalability and molecular interaction modeling. However, they face challenges in data interpretability and theoretical foundations. Current modeling approaches for DEL read count data primarily rely on theoretical prior distribution assumptions, lacking sophisticated noise handling mechanisms and robust statistical frameworks. These limitations manifest in the inability to incorporate information beyond prior distributions and insufficient reliable activity validation data for denoising, leading to suboptimal performance when addressing Distribution Noise and Distribution Shift in DEL data.
@@ 3 Method
Figure 2: Overview of DEL-Ranking framework. The model directly fuses molecule binding poses and fingerprints as input features. ARC employs target effects and binding affinity to enhance read count prediction. The ranking-based loss incorporates target effects and matrix effects for noise removal, improving the correlation between predicted read counts and true binding affinities.
@@ 3.1 Problem Formulation and preliminaries
DEL Prediction Framework. Given a DEL dataset $\mathcal{D}=\{(\mathbf{f}_{i},\mathbf{p}_{i},M_{i},R_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{f}_{i}\in\mathbb{R}^{d}$ denotes the molecular fingerprint, $\mathbf{p}_{i}\in\mathbb{R}^{m}$ represents the binding pose, $M_{i}\in\mathbb{R}$ is the matrix count derived from control experiments without protein targets, $R_{i}\in\mathbb{R}$ is the target count obtained from experiments involving protein target binding, and $y_{i}\in\{0,1\}$ indicates the activity label. We propose a joint multi-task learning framework $\mathcal{F}:\mathbb{R}^{d}\times\mathbb{R}^{m}\rightarrow\mathbb{R}\times\mathbb{R}\times[0,1]$ such that:
@@
Zero-Inflated Poisson Distribution (ZIP) & ZIP Loss. DEL screening often results in read count distributions with a high proportion of zeros due to experimental factors. To address this, previous methods (Shmilovich et al., 2023; Lim et al., 2022) have employed zero-inflated distributions, modeling read counts $r_{i}$ that can take values from either from target counts $M_{i}$ or target counts $R_{i}$. Similarly, we define $\hat{r}_{i}$in $\{\hat{M}_{i},\hat{R}_{i}\}$ as the model’s predicted read count values.
$P(X=r_{i}|\lambda,\pi)=\begin{cases}\pi+(1-\pi)e^{-\lambda},&\text{if }r_{i}=0\\ (1-\pi)\frac{\lambda^{r}_{i}e^{-\lambda}}{r_{i}!},&\text{if }r_{i}>0\end{cases}$ (2)
where $\pi$ denotes the probability of excess zeros, and $\lambda$ denotes the mean parameter of the Poisson component. In (Shmilovich et al., 2023), ZIP of $M_{i}$, $R_{i}$ are modeled respectively by different $\pi$ values, including $\pi_{M}$ and $\pi_{R}$, based on orders of magnitude, with the regression achieved by minimizing the Negative Log-Likelihood (NLL) for all predicted read counts $\hat{M}_{i}$ and $\hat{R}_{j}$.
$\mathcal{L}_{\text{ZIP}}=-\sum_{i}\log[P(\hat{M}_{i}|\lambda_{M},\pi_{M})]-\sum_{j}\log[P(\hat{R}_{j}|\lambda_{M}+\lambda_{R},\pi_{R})]$ (3)
where $\lambda_{M}$ and $\lambda_{R}$ represent Poisson mean parameters for matrix and target counts, and $\pi_{M}$ and $\pi_{R}$ denote their zero-excess probabilities. Joint modeling of target and control counts captures the relationship between DEL experiments while enhancing the model’s understanding of read count correlations with inputs.
$K_{i}$ Estimation. DEL read count prediction aims to estimate compound-target binding affinities –experimental-valided $K_{i}$ values, which are crucial for identifying promising drug candidates. We evaluate our model’s effectiveness using the Spearman rank correlation coefficient ($\rho_{s}$) between predicted read counts and true $K_{i}$ values: $\rho_{s}=1-\frac{6\sum d^{2}}{n(n^{2}-1)}$ where $n$ is the sample number and $d$ is the predicted ranking discrepancy. Ideally, $K_{i}$ values and read counts are negatively correlated.
@@ 3.2 Ranking-based Distribution Noise Removal
where $\beta\in[0,1]$ is a balancing hyperparameter to fit the relative magnitude of two components. $\mathcal{L}_{\text{PSR}}$ (Pairwise-Soft Ranking Loss) addresses local pairwise comparisons, while $\mathcal{L}_{\text{LGR}}$ (listwise Global Ranking Loss) captures global ranking information. Together, they aim to achieve a well-ordered Zero-Inflated Poisson distribution for read count values. To formally establish the effectiveness of our ranking-based approach, we provide the following theoretical justification:
@@ 3.2.1 Pairwise Soft Ranking Loss
To better model the relationships between compound pairs and handle read count noise, we introduce a novel Pairwise Soft Ranking (PSR) loss function. The PSR loss enables smooth comparison between compounds while maintaining stable optimization. $\mathcal{L}_{\text{PSR}}$ is defined as:
$\mathcal{L}_{\text{PSR}}(\hat{r}_{i},\hat{r}_{j},T)$ $=-\sum_{i=1}^{n}\hat{r}_{i}\left(\sum_{j\neq i}(\Delta_{ij}\cdot\sigma_{ij})-\sum_{j\neq i}(\Delta_{ji}\cdot\sigma_{ji})\right)$ (5)
$\sigma_{ij}$ $=\frac{1}{1+e^{-|r_{i}-r_{j}|/T}},\quad\Delta_{ij}=\frac{\Delta\text{G}_{ij}\cdot\Delta\text{D}_{ij}}{Z}$
where $\hat{r}_{i}$ and $\hat{r}_{j}$ represent the predicted read count value for compound $i$ and $j$, respectively. To ensure smooth gradients and numerical stability, we introduce a scaling factor $\sigma_{ij}$ with temperature $T$ ensures smooth transitions between rankings. To accurately model ranking changes, we design a pairwise importance term $\Delta_{ij}$ that quantifies the impact of swapping compounds $i$ and $j$. This term incorporates a gain function $G_{i}=\text{softplus}(r_{i})$ that reflects the relevance of each compound, and a rank-based discount function $D_{i}=1/(\log_{2}(\text{rank}_{i}+1)+\epsilon)$, where $\epsilon$ is a small constant for numerical stability. These gain and discount mechanisms emphasize high-ranking samples, enabling the model to focus on top-ranking cases and effectively identify potentially high-activity compounds. By combining these components into smooth Delta functions, we ensure stable gradient flow:
$\Delta G_{ij}=G_{i}-G_{j}=\text{softplus}(r_{i})-\text{softplus}(r_{j})$ (6)
$\Delta D_{ij}=D_{i}-D_{j}=\frac{1}{(\log_{2}(\text{rank}_{i}+1)+\epsilon)}-\frac{1}{(\log_{2}(\text{rank}_{j}+1)+\epsilon)}$
To normalize ranking effects, we introduce a normalization factor computed from the top-K predicted values in each batch. By choosing K smaller than the batch size N, we achieve two benefits: enhanced computational efficiency and avoidance of ranking noise from zero-value predictions.
$Z=\sum_{k=1}^{K}\frac{\text{softplus}(\hat{r}_{[k]})}{\log_{2}(k+1)+\epsilon}$ (7)
where $\hat{r}_{[k]}$ represents the k-th highest predicted read count in descending order; $\epsilon$ is set to $1e^{-8}$ to avoid division by zero. This normalization factor adjusts the loss scale across different dataset sizes and read count distributions, ensuring robust model training regardless of data variations.
@@
$\mathcal{L}_{\text{LGR}}(\hat{r},\tau,T)=-\sum_{i=1}^{N}\log\frac{\exp(\hat{r}_{\pi(i)}/{T})}{\sum_{j=i}^{N}\exp(\hat{r}_{\pi(j)}/{T})}+\sigma\sum_{i=1}^{N}\sum_{j\neq i}\mathcal{L}_{\text{con}}(\hat{r}_{i},\hat{r}_{j},\tau)$ (8)
where $\pi$ is the true ranking permutation of the compounds; $\tau$ represents the minimal margin of predicted read count pair ($r_{i}$, $r_{j}$); $T$ is a temperature parameter for score rescaling to sharpen the predicted distribution, and $\mathcal{L}_{\text{con}}$ denotes a contrastive loss among ranking scores to capture local connections, and $\sigma$ denotes the weight. This formulation is designed to achieve two critical objectives in DEL experiments: (1) Near-deterministic selection of compounds with the highest read counts, corresponding to the highest binding affinities; (2) Increased robustness to small noise perturbations in the experimental data. As $T$ approaches 0, our model becomes increasingly selective towards high-affinity compounds while maintaining resilience against common experimental noises. This dual optimization leads to more consistent identification of promising drug candidates and enhanced reliability in the face of experimental variability.
Despite the strengths of $\mathcal{L}_{\text{PSR}}$ and $\mathcal{L}_{\text{LGR}}$, they struggle to differentiate activity levels among compounds with identical read count values, particularly those affected by experimental noise. This limitation can lead to misclassification of high-activity samples with artificially low read counts as truly low-activity samples. To address this critical issue, we introduce a novel contrastive loss function $\mathcal{L}_{\text{con}}$, designed to enhance discrimination between varying levels of biological activity, especially for samples with zero or identical read count values. Let $f:\mathcal{R}\rightarrow\mathbb{R}$ be a ranking function and $\tau>0$ a fixed threshold. We define $\mathcal{L}_{\text{con}}:\mathcal{R}\times\mathcal{R}\times\mathcal{R}\rightarrow\mathbb{R}{\geq 0}$ as:
$\mathcal{L}_{\text{con}}(\hat{r}_{i},\hat{r}_{j},\tau)=\max\{0,\tau-(f(\hat{r}_{i})-f(\hat{r}_{j}))\}$ (9)
This loss function is positive if and only if $f(\hat{r}_{i})-f(\hat{r}_{j})<\tau$, enforcing a minimum margin $\tau$ between differently ranked samples. The constant gradients $\partial\mathcal{L}_{\text{con}}/\partial f(\hat{r}_{i})=-1$ and $\partial\mathcal{L}_{\text{con}}/\partial f(\hat{r}_{j})=1$ for $f(\hat{r}_{i})-f(\hat{r}_{j})<\tau$ promote robust ranking relationships.
Algorithm 1 Refinement Stage for Activity-Referenced Correction (ARC) Algorithm
1:Pose structure embeddings $\mathbf{h}_{p}$, Fingerprint sequence embeddings $\mathbf{h}_{f}$, Sequence-structure balancing weight $\varsigma$, num_iterations $\mathbf{n}$, use_feedback
2:$x\leftarrow\texttt{PostAddLayer}(\varsigma\mathbf{h}_{p}+\mathbf{h}_{f})$
3:$\hat{M}\leftarrow\texttt{MatrixHead}(\mathbf{h}_{f})$
4:Initialize $\hat{R}\leftarrow\mathbf{0}$, $\hat{y}\leftarrow\mathbf{0}$
@@
7:         $\hat{R}\leftarrow[x;\hat{p}]$, $\hat{p}\leftarrow[x;\hat{R}]$
@@
9:         $\hat{R}\leftarrow x$, $\hat{p}\leftarrow x$
@@
11:     $\hat{R}\leftarrow\texttt{EnrichmentHead}(\texttt{ReadHead}(\hat{R}))$
12:     $\hat{p}\leftarrow\texttt{ActHead}(\hat{p})$
13:end forreturn $\hat{M}$, $\hat{R}$, $\hat{p}$